OpenStudy (anonymous):

What is the solution set of this system of equations? y = x2 − 3x − 4 x = y + 8 {(-1, 0), (4, 0)} {(8, 0), (0, 8)} {(0, 0)} {(2, -6)} There is no real solution

OpenStudy (anonymous):

\[y = x^2 − 3x − 4\] \[x = y + 8\] \[x = x^2 − 3x − 4 + 8\] \[x^2 − 4x +4=0\] \[(x-2)(x-2)=0\] \[x=2 , y=x-8=2-8=-6\] \[y=-6\]